OpenStudy (anonymous):

one number is equal to the square of another. Find the numbers if both are positive and their sum is 756.

OpenStudy (anonymous):

Not quite sure!

OpenStudy (anonymous):

One number (call it x) is equal to the square of another (call it y), so x = y^2. Both are positive and their sum is 756. So x+y=756. This gives two equations. If we substitute in x = y^2 into the second equation, we get y^2 + y = 756. Subtract 756 over to get a quadratic equation: y^2 + y - 756 = 0. That factors down to (y+28)(y-27)=0, so y can either be -28 or 27, but note both x and y are positive, so y = 27. If you can't factor it, you can always use the quadratic formula. Finally, we also know x = y^2, so x=27^2=729

OpenStudy (anonymous):

pyeh is right..........

OpenStudy (anonymous):

pyeh is right..........